OpenStudy (anonymous):

-2≤4x+6<22, solve the compound inequality for x

OpenStudy (anonymous):

4x + 6 < 22 4x < 16 x < 4 4x >= -2 x >= - 1/2 -1/2 =< x < 4

OpenStudy (anonymous):

I dont understand?

OpenStudy (anonymous):

I arrived at -2<=x<4

OpenStudy (anonymous):

sorry i got second part wrong its 4x + 6 >= -2 4x >= -2 -6 = -8 x >= -2 your answer is correct for some reason i left out the +6 in 2nd part